OpenStudy (anonymous):

whats the limit of csc(2x) approaching positive pie?

OpenStudy (tkhunny):

"pie" is for eating. "pi" is for math and greek spellers. As x approaches \(\pi\) 2x approaches \(2\pi\) csc(2x) approaches \(\dfrac{1}{\sin(2x)}\) sin(2x) approaches \(0\)